OpenStudy (anonymous):

Find the arc length: (2cost, 2sint, t) for 0 <= t <= 2pi

OpenStudy (anonymous):

if you're curious, the answer is 2sqrt(5pi)

OpenStudy (anonymous):

What is this acr lenght? because the way I was thinking i got 4pi

OpenStudy (anonymous):

integrate

OpenStudy (anonymous):

this is a parametric equation intergrate from 0 to 2pi for the function given

OpenStudy (anonymous):

\[l=\int |r'(t)| \ \text{d}t \]

OpenStudy (anonymous):

c'(t) = ( -2sint, 2cost, 1) ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok, now do I use the dot product on itself?

OpenStudy (anonymous):

or sqaure each one and add then together?

OpenStudy (anonymous):

yes and finally take square root ...(thats a vector in 3D)

OpenStudy (anonymous):

all that in a square root to the integral 0 to 2pi

OpenStudy (anonymous):

i think i am messing up somewhere when i try that. can you help me with those steps?

OpenStudy (anonymous):

@mukushla use the correct term, magnitude of the vector as we hav |r'(t)|

OpenStudy (anonymous):

first find r'(t) = (-2sint, 2cost,1) magnitude of r'(t)= sqrt (4sin^2+4cos^t+1)= sqrt(5)

OpenStudy (anonymous):

then integrate sqrt(5) from 0 to 2pi

OpenStudy (anonymous):

can I show you the formula i was given?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

L(c) = ||c'(t)||dt from integral a to b, = Sqrt( c'(t) (dot prod) c'(t) ) from integral a to b

OpenStudy (anonymous):

|| || is the Norm of a vector

OpenStudy (anonymous):

which is same formula as given by @mukushla

OpenStudy (anonymous):

did not know

OpenStudy (anonymous):

look up pauls online notes calculus II or calculus III and I am sure it will be of great help to you : )